OpenStudy (anonymous):

Eliminate the arbitrary constant of 1. y=Ae^2x+Bxe^2x 2. y=x^2+C1e^2x+C2e^3x 3. y=C1e^2x+C2e^3x

OpenStudy (anonymous):

1.y=Ae^2x+Bxe^2x y'= 2Ae^2x+2Bxe^2x y''=4Ae^2x+4Bxe^2x

OpenStudy (anonymous):

.y=Aex^2x+Bxe^2x EQ. 1 y'= 2Aex^2x+2Bxe^2x EQ 2 y''=4Ae^2x+4Bxe^2x EQ 3 MULTIPLY EQ. 1,2 BY 4 4y=4Aex^2x+4Bxe^2x ------EQUATION 4 4y'=8Aex^2x + 8 Bxe^2x ---------- EQ. 5 SUBRACT EQ 4 AND 5 4y= 4Aex^2x+4Bxe^2x 4y'=8Aex^2x+8Bxe^2x 4y'-4y= -4Aex^2x-4Bxe^2x EQ 6 ADD EQ. 3 AND 6 4y'-4y=-4Aex^2x-4Bxe^2x y''=4Ae^2x+4Bxe^2x y''+4y'-4y=0 ANSWER